# AutoRev: Multi-Modal Graph Retrieval for Automated Peer-Review Generation

### v2

@@ 4.2.2. Edge Definitions
4.3. GNN Training for Multi Modal Passage Retrieval
Figure 3. Excerpt of a test sample where passages highlighted in cyan colour refer to relevant passages identified by GNN only and yellow colour refer to relevant passages identified by both GNN and DPR:
To train a GNN on the graph constructed in Section 4.2, we take the data generated in Section 4.1, which maps each paper $p_{i}$ to a set of important passages, $\mathcal{O}_{i}=\{o_{1},o_{2},\dots,o_{n}\}$. Each passage is important for generating a review for the paper it is associated with. This mapping is used as a training set for training the GNN. The node embeddings for sentence nodes are initialized with embeddings from a Sentence-Transformer (Reimers and Gurevych, 2019) model, namely all-mpnet-base-v2. The node embeddings for all other nodes are initialized with zero embeddings of dimension $768$, matching that of the Sentence-Transformer model.
Now, a Graph Attention Network (Veličković et al., 2018) is used to train the graph to predict the important passages. After passing the initial representation of the graph nodes through three layers of GATConv (PyTorch Geometric’s implementation of Graph Attention Networks) [^6] https://pytorch-geometric.readthedocs.io/en/latest/generated/torch_geometric.nn.conv.GATConv.html, we apply a non-linear classifier to the final passage node representations. The classification is defined as follows:
  1. (1)
    A label of 0 indicates that the passage is not important for review generation.
  2. (2)
    A label of 1 indicates that the passage is important for review generation.
The GNN is trained using various configurations based on the values of $k$ and $m$, as described in the DPR process in Section 4.1. These trained GNN models are denoted as GNN $(k,m)$.
We hypothesize that our choice of graph construction which models the semantic representation of nodes and flow of information through structure, allows our model to capture the intricate inter-dependencies of sentences and passages across different sections within an academic paper, and effectively make a choice about which passage(s) is/are important. The dependencies, according to our hypothesis, are learned from the sequential and hierarchical connections that we introduce in our graph construction method. The passing of information from one node to another is enhanced by the GATConv (Equations 2, 3, and 4) layers, which involves the concept of attention to selectively propagate messages in the Message Passing layer of a GNN.
(2) $\mathbf{h}_{i}^{\prime}=\sigma\left(\sum_{j\in\mathcal{N}(i)}\alpha_{ij}\mathbf{W}\mathbf{h}_{j}\right)$
(3) $e_{ij}=\text{LeakyReLU}\Bigl(\mathbf{a}^{T}\bigl[\mathbf{W}\mathbf{h}_{i}\parallel\mathbf{W}\mathbf{h}_{j}\bigr]\Bigr)$
(4) $\alpha_{ij}=\frac{\exp\bigl(e_{ij}\bigr)}{\sum_{k\in\mathcal{N}(i)}\exp\bigl(e_{ik}\bigr)}$
Additional implementation details related to this section are available in Appendix B.
As illustrated in the example from our test set (Figure 3), the DPR-based approach fails to retrieve a key passage related to the main contributions of the paper, whereas our GNN successfully identifies it. Therefore, our multi-modal graph, which leverages both sequential and hierarchical node connections, more effectively identifies the salient passages required to enhance the review generation process. The GNN extracted passages, along with a structured prompt, serve as the input in the LLM fine-tuning stage, as discussed in Section 4.5.
4.4. Grounding Requirements and Inference Efficiency
For practical deployment, our GNN-based method offers a crucial advantage by operating autonomously at inference time. It identifies important passages directly from the paper’s graph structure, eliminating the need for grounding from human-generated reviews which act as queries in the DPR-based approaches, but may be unavailable in real-world scenarios.
Second, by selecting only the passages deemed most relevant, we significantly reduce the input length and associated computational costs for downstream LLM-based review generation. This not only leads to higher performance of LLMs, as outlined by Hsieh et al. (2024) but also leads to higher scores of the generated reviews compared to our baselines (see Section 5). The statistics related to the decrease in number of passages and token length post-extraction by the GNN model are reported in Table 2.
Table 2. Average token length (Avg. Token Len) and Average passage count (Avg. Passages) for a full paper, our best-configuration: GNN $(5,5)$, and our fastest-configuration: GNN $(5,3)$.
Input to LLM Avg Token Len Avg Passages
Full Paper 9445.92 47.07
GNN $(5,5)$ 4147.15 29.49
GNN $(5,3)$ 2899.59 21.00
4.5. Fine-Tuning Large Language Models for Review Generation
General-purpose LLMs are not trained specifically to generate peer reviews for academic research papers. Our motivation for fine-tuning is to ensure that the models learn to generate qualitatively sound, factual and structured peer reviews.
We select Mistral-Instruct-v0.2 (Mistral) [^7] https://arxiv.org/abs/2310.06825, inspired by our baseline SEA-EA, which also uses Mistral-Instruct-v0.2 as its underlying backbone model. We also use Llama 3.1 8B (Llama) [^8] https://arxiv.org/abs/2407.21783, as it is an open-weight state-of-the-art (SOTA) model with strong text generation capabilities for a variety of tasks. We employ LoRA (Low-Rank Adaptation) (Hu et al., 2022), for fine-tuning due to its efficiency in adapting large language models to particular tasks while maintaining computational feasibility. This allows us to leverage the pre-trained capabilities of Llama and Mistral while efficiently adapting them for review generation.
We use the trained GNN, as described in Section 4.3, to retrieve relevant passages for each paper that serve as the LLM’s input during fine-tuning. The human-written peer reviews for ICLR 2024 papers, which are consolidated with GPT-4, served as the ground truth for fine-tuning the LLM used to generate the final reviews. These reviews are released as a part of the dataset released by Yu et al. (2024). The dataset split is summarized in Table 3.
Table 3. Number of samples used for training and testing sourced from Yu et al. (2024)
Dataset Split Number of Samples
Training Set 1,765
Test Set 271
Both Mistral and Llama are fine-tuned on the passages retrieved, using various configurations of the trained GNN $(k,m)$, as described in Section 4.3. These fine-tuned models are denoted Mistral $(k,m)$ and Llama $(k,m)$, respectively. The prompt and hyperparameter settings used during fine-tuning are provided in Appendix C.
5. Results and Analysis
Table 4. Overall performance on the ICLR-24 reviews test dataset, comparing M-7B, SEA-E, SEA-EA as baselines from Yu et al. (2024) with our fine-tuned models. For both Llama-3.1-8B (Llama) and Mistral-Instruct-v0.2 (Mistral), the configuration with $(5,5)$ achieves the best performance, while $(5,3)$ has the shortest training time. Bold highlights the best scores, and underlined values indicate the second-best. Please note that the GPU configurations for finetuning Llama and Mistral were different. For details, refer to Appendix C.
Model ROUGE F1 ROUGE Recall BERTScore-F1 Training Time (hrs)
R-1 R-2 R-L R-1 R-2 R-L
M-7B 22.55 8.89 11.28 13.48 5.29 16.73 83.79
SEA-E 34.29 11.71 14.98 23.28 7.90 10.13 84.04
SEA-EA 36.97 13.02 15.88 25.62 8.97 10.97 84.15
Mistral $(5,3)$ 53.44 15.47 22.39 53.18 15.42 22.31 86.44 4.34
Llama $(5,3)$ 54.24 15.91 22.59 53.75 15.77 22.41 86.59 3.60
Mistral $(5,5)$ 53.65 15.80 22.74 53.60 15.79 22.74 86.55 6.44
Llama $(5,5)$ 54.47 16.26 22.87 53.49 15.98 22.48 86.68 5.30
This section details our system’s performance analysis. Section 5.1 reports the evaluation of generated review quality using ROUGE-1, ROUGE-2, and ROUGE-L (Recall and F1) (Lin, 2004), as well as BERTScore-F1 (Zhang* et al., 2020), on the GPT-4-consolidated reviews test set for ICLR 2024 papers. ROUGE measures n-gram overlap, with recall reflecting content coverage and F1 balancing precision and recall, while BERTScore captures semantic similarity in embedding space. Section 5.2 further employs the Aspect Coverage (ACov) metric (Yuan et al., 2022), designed specifically for assessing peer-review quality.
Given the limitations of automated metrics, we also conduct human evaluations (Section 5.3) to assess review acceptability. To address the inherent scalability challenges of manual assessment, we employ an LLM-as-a-Judge framework (Section 5.4), which shows strong correlation with human judgments.
5.1. Performance Comparison with Baselines
We compare the performance of our models against baselines sourced from Yu et al. (2024). As reported in Table 4, the M-7B baseline performs direct inference with Mistral guided by an instruction prompt to generate peer reviews, SEA-E baseline leverages standardized fine-tuning data to produce more constructive reviews, while the SEA-EA baseline further introduces a mismatch score to evaluate the consistency between paper content and generated reviews. Our results suggest that we consistently outperform the baselines on all metrics, with significant improvements in the ROUGE metrics (Recall and F1). Our best-performing configuration, Llama $(5,5)$ (selected based on the highest BERTScore), achieves an improvement of 58.72% over the state-of-the-art (SOTA) baselines across all evaluation metrics. To ensure that these gains are not solely due to using a more capable LLM, we also evaluate our method using Mistral. The results demonstrate that even with Mistral, integrating our Multi Modal GNN-based passage selection method leads to substantial improvements, confirming that the performance boost is primarily due to our graph-based retrieval strategy rather than the choice of the LLM alone. The integration of a Graph Attention Network (GAT) enables the model to effectively capture complex dependencies between relevant passages and sentences within the paper, offering the essential context required to generate comprehensive reviews.
Our fastest, second-best, fine-tuned model, Llama $(5,3)$, performs competitively while being highly efficient. It reduces the training time by 32.07% with only a 0.71% drop in performance compared to our best model, consistently outperforming all baselines. This shows our GNN-based approach achieves near-optimal performance with significantly less computational overhead.
5.2. Aspect Coverage Analysis
As discussed by Yuan et al. (2022), each comment in a peer review typically addresses diverse aspects of an academic research paper, such as Summary (SUM), Motivation/Impact (MOT), Originality (ORI), Soundness/Correctness (SOU), Substance (SUB), Replicability (REP), Meaningful Comparison (CMP), and Clarity (CLA). These cannot be reliably assessed by conventional automatic metrics like ROUGE-F1 or BERTScore. To address this, Yuan et al. (2022) introduces an Aspect Tagger which can be used to compute Aspect Coverage (ACov) that measures the proportion of aspects covered by a peer-review. We evaluate the two best-performing models Llama $(5,5)$ and Llama $(5,3)$ using this tagger.
Table 5. Aspect Coverage (%) of model-generated reviews evaluated using the Aspect Tagger (Yuan et al., 2022). Summary (SUM), Clarity (CLA), Soundness/Correctness (SOU), Substance (SUB), Originality (ORI), Meaningful Comparison (CMP), Replicability (REP), and Motivation/Impact (MOT).
Model SUM CLA SOU SUB ORI CMP REP MOT
Llama (5,3) 100.0 99.6 98.9 95.2 92.6 79.0 72.0 66.8
Llama (5,5) 100.0 99.3 99.3 95.9 90.0 84.5 71.6 66.8
As shown in Table 5, both variants achieve strong coverage for Summary, Clarity, and Soundness, indicating the reliability of our model-generated reviews in capturing these core aspects. However, the aspect coverage for Motivation/Impact and Replicability is comparatively lower. Overall, the results suggest that our model captures most key aspects, indicating that the generated reviews are comprehensive and address the quality of multiple dimensions of a research paper.
5.3. Human Evaluation
Understanding the limitations of automatic evaluation metrics as highlighted by Hanna and Bojar (2021) and Schluter (2017), we conduct human evaluations to assess the acceptability of the generated reviews. We model our evaluation around two distinct aspects of review quality. The first is a Qualitative Assessment, where the generated review is evaluated in isolation across four key criteria: Reviewer Confidence, Review Thoroughness, Review Constructiveness, and Review Helpfulness. The second is a Factual Validity Assessment, which ensures the review’s specific claims are factually grounded in the source document. Since human evaluation is very costly, we conduct each of the two tasks on a subset of the test dataset, as described in their individual subsections. Additionally, we only considered Llama $(5,3)$ and Llama $(5,5)$, our best performing models for the evaluation process.
5.3.1. Qualitative Assessment
A total of 30 papers are chosen at random from our test set. Each of these papers are associated with 3 reviews, one Gold, one from Llama $(5,3)$, and one from Llama $(5,5)$, thus, resulting in a total set of 90 reviews. A total of 27 annotators are chosen, and each annotator is given 10 reviews to annotate. The evaluation criteria is designed to capture key aspects of review quality and include:
  1. (1)
    Reviewer Confidence (RCf): Reflects the reviews’ depth of technical understanding and engagement with the content.
  2. (2)
    Review Thoroughness (RT): Measures how well the review covers the methodology, results, and reproducibility.
  3. (3)
    Review Constructiveness (RCt): Assesses the clarity and usefulness of suggestions for improvement.
  4. (4)
    Review Helpfulness (RH): Indicates how well the review identifies important issues and explains them clearly to the authors.
Each annotator is tasked to assign a score from 1 to 5 for each of the above criteria. The evaluation guidelines are detailed in Appendix D. We end up with each review being annotated by 3 annotators. Score for a review is averaged across all its annotators, and the final score of a model, is the average score across all its reviews. The results of this experiment are summarized in Table 6.
Table 6. Results of our human qualitative assessment. This table shows average scores (on a scale of 1-5) comparing gold-standard reviews with those generated by our two best-performing Llama models across four key criteria: Reviewer Confidence (RCf), Review Thoroughness (RT), Review Constructiveness (RCt), and Review Helpfulness (RH).
Review RCf RT RCt RH
Gold 3.54 3.49 3.26 3.41
Llama (5,3) 3.37 3.30 3.27 3.34
Llama (5,5) 3.27 3.24 3.19 3.30
5.3.2. Factual Validity Assessment
A total of 10 papers are chosen at random from our test set. Each of these papers are associated with 3 reviews, one Gold, and two generated by our best performing models. Thus, the final evaluation set consists of 10 papers, each associated with 3 reviews, one Gold, one from Llama $(5,3)$, and one from Llama $(5,5)$. These 10 papers are circulated among 10 annotators. It is ensured that a single paper is annotated by at least 2 annotators. Annotators are instructed to evaluate each statement within a review based on its factual grounding in the full source paper. For each point, they assign a rating from a four-point scale: ‘Completely Agree’, ‘Mostly Agree’, ‘Mostly Disagree’, or ‘Completely Disagree’. The evaluation guidelines are detailed in Appendix D. The results of this experiment are summarized in Table 7 and Table 8.
Table 7. Detailed results of the Factual Validity Assessment. This table presents the percentage distribution of human agreement ratings for claims made in reviews from our Llama models and the gold standard. The abbreviations are: Completely Agree (CA), Mostly Agree (MA), Mostly Disagree (MD), and Completely Disagree (CD).
Review CA MA MD CD
Gold 39.12% 37.85% 16.40% 6.62%
Llama (5,3) 32.64% 31.94% 22.92% 12.50%
Llama (5,5) 32.27% 32.98% 23.05% 11.70%
Table 8. Detailed results of the Factual Validity Assessment. This table shows the overall Positive Rating Percentage, calculated by summing the “Completely Agree” and “Mostly Agree” categories from Table 7, which indicates the factual soundness of the reviews.
Review Positive Rating %
Gold 76.97%
Llama (5,3) 65.25%
Llama (5,5) 64.58%
5.4. LLM-as-a-Judge Evaluation
As mentioned by Zheng et al. (2023), the LLM-as-a-Judge method offers an explainable and affordable way to approximate human preferences, which are otherwise costly to gather. We employ a large language model, Gemini-2.0-Flash [^9] https://deepmind.google/technologies/gemini/ as an automated judge to evaluate the quality of reviews generated by our proposed model as compared to the gold-standard reviews. For the LLM-as-a-judge task, the evaluation criteria and prompting guidelines are adapted from those used in our human qualitative assessment, which are detailed in Section 5.3.1. Since LLM-as-a-Judge evaluation is scalable and much cheaper to undertake, we evaluate all our models against the gold review, over the entire test dataset. A key finding from our evaluation is the strong correlation between the LLM-as-a-Judge scores (Table 9) and the results from our human annotators (Table 6). This alignment suggests that the LLM-as-a-Judge is a reliable and valid proxy for assessing review quality in this domain.
Table 9. Qualitative evaluation of model-generated reviews using Gemini-2.0-Flash as the LLM judge. The table presents the average scores (out of 5) for all model configurations and the gold standard across four criteria: Reviewer Confidence (RCf), Review Thoroughness (RT), Review Constructiveness (RCt), and Review Helpfulness (RH). Bold values indicate the best score in each column (excluding Gold), and underlined values indicate the second-best.
Review RCf RT RCt RH
Gold 3.98 3.87 3.79 3.79
Mistral $(5,3)$ 3.50 3.35 3.44 3.32
Llama $(5,3)$ 3.61 3.45 3.47 3.39
Mistral $(5,5)$ 3.48 3.36 3.43 3.34
Llama $(5,5)$ 3.63 3.44 3.47 3.41
5.5. Ablation Studies
We perform ablation studies by systematically varying the values of $k\in\{1,3,5\}$ and $m\in\{3,5\}$ during LLM fine-tuning. The detailed results are reported in Appendix A. The results indicate that increasing $k$ and $m$ provides richer context to both Mistral and Llama to generate reviews, leading to consistent improvements across all evaluation metrics. Additionally, we observe that higher values of $m$ result in longer training times, as the GNN retrieves and passes a larger set of passages as input for finetuning the LLMs.
6. Limitations
While our framework provides a robust and high-performing solution for review generation, we acknowledge that no automatic system can fully replicate human judgment in complex tasks such as peer reviewing, as noted by Ye et al. (2024). Our goal is to assist authors in refining their papers and we aim for our system to support human reviewers. Currently, our system is primarily focused on ICLR 2024 papers due to the limited availability of public datasets containing consolidated reviews. However, we believe in the generalization capabilities of our model, given the proven effectiveness of graph-based approaches in various downstream NLP tasks, as discussed in Section 1.
Our reliance on human written reviews, consolidated using GPT-4 represents another limitation. While multi-review datasets more closely reflect the diversity of real peer review, they are often noisy and contradictory, which can hinder model learning. Consolidated reviews provide a clearer and more consistent supervision signal, but at the cost of reducing the richness of differing reviewer perspectives. Future work should explore hybrid approaches that retain both consistency and diversity.
7. Ethical Considerations
The deployment of AutoRev raises key ethical concerns related to bias, misuse, and academic integrity. Since the system is trained on existing peer reviews, it may inherit and amplify biases present in those texts, such as leniency toward well-known authors or under-representation of certain research areas, potentially skewing evaluations. Moreover, there is a risk that users may over-rely on AutoRev for making or substituting final review decisions, which could undermine the depth and critical thinking central to human peer review. Additionally, the realistic quality of AI-generated reviews opens the possibility of unethical use, such as submitting AI-authored critiques as original human work. To address these issues, we advocate for responsible usage of AutoRev as a supplementary tool, accompanied by transparency in AI involvement and safeguards to preserve academic integrity.
8. Conclusion
In this work, we introduce AutoRev, an automatic peer review generation system designed specifically for academic research papers. Our approach leverages a novel graph-based framework to represent the structure of a paper which enables the effective extraction of important passages using a Graph Neural Network. This strategy of selecting passages and reducing the input length to LLMs addresses the limitations of traditional methods, like finetuning using the entire paper as input (Yu et al., 2024), which leads to sub-optimal performance of LLMs due to large input token length (Hsieh et al., 2024).
Unlike DPR-based methods that require initial grounding on human written references to retrieve relevant passages, AutoRev employs a GNN that learns to identify important content directly from the paper’s graphical representation, without the need for grounding on existing reviews. This not only streamlines the process but also reduces computational overhead while enhancing review quality.
Our experiments on the ICLR 2024 dataset demonstrate that AutoRev consistently outperforms existing baselines, by 58.72% on popular evaluation metrics like ROUGE, and BERTScore. These findings are further strongly backed by human evaluations, where the generated reviews were rated nearly as constructive and acceptable as human-written ones, suggesting AutoRev could meaningfully reduce reviewer workload.
Beyond these performance gains, we emphasize AutoRev’s broader role in the web ecosystem. Academic research papers represent some of the most authoritative and trusted content on the web. By accelerating the peer-review workflow and enabling scalable, automated feedback, AutoRev contributes to the timely dissemination of credible scholarly knowledge online. In doing so, it supports not only authors and reviewers but also the wider web community that depends on accurate, trustworthy research for discovery, indexing, and downstream applications. Future work will explore extending this framework to a broader range of academic domains and investigating further improvements in graph construction and neural retrieval techniques to continue advancing the state-of-the-art in automated review generation.
Appendix A Ablation studies
The quantitative performance of different combinations of $k$ (number of review sentences concatenated; $k=1,3,5$) and $m$ (number of top DPR-retrieved passages used to train the GNN; $m=3,5$) for review generation is reported in Table 10 and Table 11, evaluated against the gold-standard reviews. While calculating the ROUGE metrics, we enable the use_stemmer option to ensure consistent matching of words with different morphological forms, to improve the reliability of the evaluation.
Table 10. Overall performance of Mistral across different $(k,m)$ settings. Metrics include ROUGE-F1 (R-1, R-2, R-L), ROUGE-Recall (R-1, R-2, R-L), BERTScore F1 and training time (in hours). Bold highlights the best scores, and underlined values indicate the second-best.
@@ Appendix A Ablation studies
Table 11. Overall performance of Llama across different $(k,m)$ settings. Metrics include ROUGE-F1 (R-1, R-2, R-L), ROUGE-Recall (R-1, R-2, R-L), BERTScore F1 and training time (in hours). Bold highlights the best scores, and underlined values indicate the second-best.
@@
Appendix B GNN Implementation Details
@@ B.2. Training
Appendix C Finetuning Implementation Details
@@
C.1. Prompt for Finetuning
The prompt used for fine-tuning both Llama 3.1 8B and Mistral-Instruct-v0.2 is shown in Listing C.1.
@@ C.1. Prompt for Finetuning
The extracted relevant passages are inserted ahead of the ”Research Paper Passages” heading. Similarly, the corresponding content for each review section (Summary, Strengths, Weaknesses, and Questions) is added after their respective headings in the generated review.
Appendix D Human Evaluation
Due to space constraints, the detailed evaluation guidelines are made available in an anonymized GitHub repository: Anonymous Repository. For the camera-ready version, we will release these materials in a public GitHub repository to ensure full transparency and accessibility. We will now proceed to show the snapshots of our annotation UI, developed using Streamlit  [^10] https://streamlit.io/.
D.1. Qualitative Assessment

### v3

@@ A.2 Edge Definitions
Appendix B Ablation studies
The quantitative performance of different combinations of $k$ (number of review sentences concatenated; $k=1,3,5$) and $m$ (number of top DPR-retrieved passages used to train the GNN; $m=3,5$) for review generation is reported in Table 9 and Table 10, evaluated against the reference feedback. While calculating the ROUGE metrics, we enable the use_stemmer option to ensure consistent matching of words with different morphological forms, to improve the reliability of the evaluation.
@@ Appendix B Ablation studies
Table 9: Overall performance of Mistral across different $(k,m)$ settings. Metrics include ROUGE-F1 (R-1, R-2, R-L), ROUGE-Recall (R-1, R-2, R-L), BERTScore F1 and training time (in hours). Bold highlights the best scores, and underlined values indicate the second-best.
@@
Table 10: Overall performance of Llama across different $(k,m)$ settings. Metrics include ROUGE-F1 (R-1, R-2, R-L), ROUGE-Recall (R-1, R-2, R-L), BERTScore F1 and training time (in hours). Bold highlights the best scores, and underlined values indicate the second-best.
Appendix C GNN Implementation Details
The passing of information from one node to another is enhanced by the GATConv (Equations 2, 3, and 4) layers, which involves the concept of attention to selectively propagate messages in the Message Passing layer of a GNN.
$\mathbf{h}_{i}^{\prime}=\sigma\left(\sum_{j\in\mathcal{N}(i)}\alpha_{ij}\mathbf{W}\mathbf{h}_{j}\right)$ (2)
$e_{ij}=\text{LeakyReLU}\Bigl(\mathbf{a}^{T}\bigl[\mathbf{W}\mathbf{h}_{i}\parallel\mathbf{W}\mathbf{h}_{j}\bigr]\Bigr)$ (3)
$\alpha_{ij}=\frac{\exp\bigl(e_{ij}\bigr)}{\sum_{k\in\mathcal{N}(i)}\exp\bigl(e_{ik}\bigr)}$ (4)
@@ C.2 Training
Appendix D Finetuning Implementation Details
@@
Table 11: Hyperparameters Used During Finetuning
D.1 Prompt for Finetuning
The prompt used for fine-tuning both Llama 3.1 8B and Mistral-Instruct-v0.2 is shown in Listing D.1.
@@ D.1 Prompt for Finetuning
The extracted relevant passages are inserted ahead of the "Research Paper Passages" heading. Similarly, the corresponding content for each feedback section (Summary, Strengths, Weaknesses, and Questions) is added after their respective headings in the generated feedback.
Appendix E LLM-as-a-judge and Qualitative Assessment
The prompt used for LLM-as-a-judge with Gemini-2.0-Flash is shown in Figure 4. The corresponding model-generated feedback is inserted under the appropriate subheadings within the prompt.
For Qualitative Assessment, the prompt in Figure 4 is modified only slightly as to just remove the output format. This slightly modified prompt is given as the evaluation guideline to the annotators. It is ensured that the annotators selected are eligible to review papers for the particular conference.
** Instruction ** : You are an expert on research papers. Evaluate the following feedback according to the scoring
[…]
Give rating for the categories described below on the full scale from 1(lowest rating) to 5(highest rating).
Assign scores independently per category (they may differ significantly).
**Category 1: Feedback Depth**
1 – Unqualified: Admits lack of expertise | Fundamental misunderstandings | Fails to engage with technical content
2 – Limited: Surface-level comprehension | Misses >2 key concepts | Confuses established methods
3 – Functional: Grasps main contributions | Identifies 1-2 technical flaws | Limited literature context
4 – Qualified: Precise technical critique | Contextualizes in recent work (>3 citations) |
Identifies methodological tradeoffs
5 – Authoritative: Demonstrates field leadership | Predicts follow-up research directions |
Reveals fundamental limitations in approach
**Category 2: Feedback Thoroughness**
1 – Negligent: <3 substantive comments | Skips methodology/results | Fails to address claims
2 – Patchy: Only addresses obvious aspects | 0 analysis of experiments | No ethical consideration
3 – Standard: Checks main claims | Basic methodology feedback | 1 reproducibility concern noted
4 – Rigorous: Line-by-line methodology check | Statistical validity analysis | 2+ improvement suggestions
5 – Forensic: Alternative approach proposals | Re-analysis of key results |
Compliance checklist (ethics, reproducibility, etc.)
**Category 3: Feedback Constructiveness**
1 – Destructive: Personal attacks | Unsubstantiated dismissal | 0 actionable feedback
2 – Demanding: Unrealistic requests ("redo all experiments") | Vague quality complaints
3 – Directive: General improvement areas | 1-2 specific examples | Mixed positive/negative tone
4 – Strategic: Priority-ranked suggestions | Alternative methodology paths | Conference-specific improvements
5 – Transformative: Step-by-step revision roadmap | Code/experiment snippets provided | Before-after examples
**Category 4: Feedback Helpfulness**
1 – Misleading: Factually incorrect suggestions | Contradicts paper’s goals | Harmful recommendations
2 – Abstract: "Improve writing" without examples | "Add experiments" without design
3 – Guided: Section-specific edits | 2-3 implementable suggestions | Partial error diagnosis
4 – Impactful: Directly address weaknesses | Templates/examples provided | Clear priority levels
5 – Revelatory: Exposes fundamental flaws | Provides benchmark comparisons | Supplies missing literature
**Scoring Methodology**
For EACH category:
1. Find the HIGHEST level (5-1) where ALL criteria are FULLY satisfied (Base Level)
2. Check criteria in the NEXT HIGHER LEVEL (Base Level + 1):
   - Calculate partial credit: (Satisfied Criteria in Higher Level / Total Criteria in Higher Level)
3. Final Score = Base Level + Partial Credit
4. Round to 2 decimal places (0.33 increments for 3-criteria levels, 0.5 for 2-criteria levels)
[…]
Respond ONLY with this JSON format:
  "feedback_scores": {{
    "Category 1": "",
    "Category 2": "",
    "Category 3": "",
    "Category 4": ""
Figure 4: Prompt used during evaluating the quality of generated feedbacks with LLM as a Judge
Appendix F Human Evaluation UI
Our annotation UI is developed using Streamlit  [^4] https://streamlit.io/.
F.1 Qualitative Assessment